A car travels 1/4 of a circle of radius r .Ratio between distance to its displacement is
1
Expert's answer
2018-09-12T09:39:07-0400
The distance is
d=2πr/4=πr/2.
The displacements is
s=√2 r.
Ratio between distance to its displacement is
d/s=(πr/2)/(√2 r)=π/(2√2).